A future contract exist on a stock with a current price of sh 80. The future contract
natures in 6 months. The risk free rate is 7% per year and the stock has an
annual dividend yield of 3%. Calculate the future price.
1
Expert's answer
2019-02-21T10:48:08-0500

S = sh 80, t =6 months, Rf = 7% per year, q = 3%.

The future price is:

F = S*e^((Rf - q)*t) = 80*e^((0.07 - 0.03)*0.5) = sh 81.62.